Consider the sequence 3,9,15,__,57 and describe how you get the 4rd term after 15 and write a general formula given tn=_____
Please help out on how to do this I'm not sure which squence it is since its a grade 9 assignment

Answer by rothauserc(4718)

3,9,15,__,57
This is an arithmetic sequence defined by the formula for the nth term in the sequence as
Xn = a + d(n - 1) where a is the first term and d is the difference
in our example a = 3 and d = 6 ( 9 - 3)
now for the fourth term
X4 = 3 + 6*(4 - 1)
X4 = 3 + 18
X4 = 21